# Adapting Large Language Models for Document-Level Machine Translation

### v2

@@ 7 Conclusion
This study investigates the adaptation of large language models (LLMs) for document-level machine translation (DocMT) through extensive experimentation with two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Results demonstrate that task-specific supervised fine-tuning on parallel documents significantly boosts the performance of moderately-sized LLM-based models (with $7B$ parameters) in DocMT, surpassing gpt-4-turbo in some cases. Our analysis offers insights into LLM-based DocMT models, providing a foundation for future advancements in the field of DocMT.
@@ Appendix B Optimization and Hyperparameters
When fine-tuning L-7B-LoRA and V-7B-LoRA on parallel documents, we employ a learning rate of $5\times 10^{-5}$ and utilize a batch size of $64$. Additionally, we apply a linear learning rate schedule, with a warm-up phase comprising $10\%$ of the total training steps. The LoRA rank is set to $16$, impacting only $0.1\%$ of the parameters (about 8M parameters). We maintain the same hyperparameters for fine-tuning Doc2Doc-mT5 models, with the exception of using a learning rate of $5\times 10^{-4}$. In this phase, L-7B-LoRA and V-7B-LoRA are fine-tuned for a maximum of $3$ epochs, and Doc2Doc-mT5 models are fine-tuned for a maximum of $10$ epochs. Early stopping is applied on the validation loss.
@@ Appendix F Off-Target Translation
1204060801000020202020404040406060606080808080100100100100Percentage (%) of training dataCOMETEnglish-Romanian
1204060801000020202020404040406060606080808080100100100100Percentage (%) of training dataCOMETEnglish-Chinese
@@
Appendix G Scaling Law of Parallel Documents from English to Romanian and Chinese

### v3

@@ 7 Conclusion
This study investigates the adaptation of LLMs for DocMT through extensive experimentation. The research involves two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Our findings indicate that while specialized models sometimes surpass GPT-4 in translation accuracy, they continue to face challenges such as off-target translation due to error propagation during decoding. We present a detailed analysis of these LLMs tailored for DocMT, exploring aspects such as translation errors, discourse phenomena, training strategies, the scaling law of parallel documents, recent test set evaluations, and zero-shot crosslingual transfer. The results highlight both the strengths and limitations of LLM-based DocMT models, offering a foundation for future research in this field.
@@ Appendix B Optimization and Hyperparameters
When fine-tuning L-7B-LoRA and V-7B-LoRA on parallel documents, we use a learning rate of $5\times 10^{-5}$ and a batch size of $64$. We apply a linear learning rate schedule with a 10% warm-up phase. The LoRA rank is set to 16, affecting only 0.1% of the parameters (about 8M parameters). The same hyperparameters are used for fine-tuning Doc2Doc-mT5 models, except for a learning rate of $5\times 10^{-4}$. L-7B-LoRA and V-7B-LoRA are fine-tuned for up to $3$ epochs, and Doc2Doc-mT5 models for up to $10$ epochs. Early stopping is based on validation loss.
@@ Appendix F Off-Target Translation
En-De En-Fr
Doc2Doc-mT5-1.2B 39.8 26.2
L-7B-LoRA 64.4 29.9
L-7B-fft 62.4 28.4
B-7B-LoRA 58.3 23.6
B-7B-fft 49.8 25.1
V-7B-LoRA 48.9 30.2
V-7B-fft 43.5 27.2
Table 18: Generative accuracy (in %) on the English-German and English-French contrastive test sets. Best results are highlighted in bold.
Appendix G Discourse Phenomena
Post and Junczys-Dowmunt (2023) propose to evaluate the accuracy on the constrastive test sets in a generative way. Hence, we present the generative accuracy in Table 18.
1204060801000020202020404040406060606080808080100100100100Percentage (%) of training dataCOMETEnglish-RomanianL-7B-LoRAL-7B-fftB-7B-LoRAB-7B-fftV-7B-LoRAV-7B-fft
1204060801000020202020404040406060606080808080100100100100Percentage (%) of training dataCOMETEnglish-ChineseL-7B-LoRAL-7B-fftB-7B-LoRAB-7B-fftV-7B-LoRAV-7B-fft
@@ Appendix G Discourse Phenomena
Appendix H Scaling Law of Parallel Documents from English to Romanian and Chinese
@@
Appendix I Zero-Shot Crosslingual Transfer
$\mu_{\Delta}$ Ar De Fr It Ja Ko Nl Ro Zh
L-7B-LoRA +29.4 +36.3 +38.8 +37.2 +32.1 +15.9 +17.1 +21.7 +35.8 +29.5
L-7B-fft +29.0 +41.2 +40.5 +37.1 +18.0 +27.7 +29.4 +11.2 +18.5 +37.5
B-7B-LoRA +20.3 0+7.5 +40.7 +20.7 +21.9 +17.5 +15.9 +23.7 +25.3 0+9.8
B-7B-fft +27.3 +14.8 +37.8 +28.9 +43.3 +13.1 +15.3 +38.5 +34.7 +19.5
We present the complete results on zero-shot crosslingual transfer in Table 19.